Question 6 (1 point) Which of the following sets of scores has the greatest variability (range)? 1) 2, 5, 8, 11 2) 13, 13, 13, 13 3) 20, 25, 26, 27 4) 42,43, 44,45 Question 7 (1 point) What is the mean of the following scores? 2,5,4,1,8 1) 3 2) 4 3) 5 4) 20
Answer #1

Q6. To see which set of score have the greatest variability(range), we have to check all four options.

a) range=11-2 = 9

b)range=13-13 =0

c) range=27-20 =7

d)range=45-42 =3

So, set of scores of option a have the highest variability.

Q7. Mean of the scores 2,5,4,1, 8 is

Mean=(2+5+4+1+8)/5

=20/5 = 4

So the correct answer is option b .